OpenStudy (anonymous):

Worded linear equation. Can't get my head around what the equation could be. A gem of weight 31 carats was broken in 2 pieces. The 2 pieces were placed on a balance, one on either side. An additional 5 carats were needed to achieve equalibrium. Determine the weight, in carats, of the larger piece.

OpenStudy (anonymous):

x - one piece y - another piece Total sum of pieces is 31 One piece weight 5 carats less so we need to add 5 to one of them \[\begin{cases}x+y=31\\x=y+5\end{cases}\] I hope you can solve it now yourself

OpenStudy (anonymous):

Oh okay... all the work I have done so far has only had one variable. But thanks, I think I know how to solve this, put y + 5 in the place of x... Thankyou very much.

OpenStudy (anonymous):

yes then you get y and later put y into second one

OpenStudy (anonymous):

cool thanks :)